About 18% of the population is allergic to pollen. If two people are randomly selected, what is the probability both are allergic to pollen? What is the probability at least one is allergic to pollen?
a)The probability that both will be allergic to pollen is?
b)The probability that at least one is allergic to pollen?

Answer by stanbon(75887)

About 18% of the population is allergic to pollen. If two people are randomly selected, what is the probability both are allergic to pollen? What is the probability at least one is allergic to pollen?
a)The probability that both will be allergic to pollen is?
Ans: 0.18^2 = 0.0324
-----------------------

b)The probability that at least one is allergic to pollen?
P(at least one allergic) = 1 - P(neither is allergic)
= 1 - (0.82)^2
----
= 1 - 0.6724
---
= 0.3276
=================
cheers,
Stan H.